With M=mass of the disk, m= hanging mass, R=radius of the disk, r=radius of the spool , a= linear acceleration , and α=angular acceleration , which formula gives the torque of the tension force, τ?

  • A. τ=mgr
  • B. τ=mαR
  • C. τ=m(g-a)r
  • D. τ=m(g-a)R
  • E. τ=(M-m)(g-a)R
  • F. None of the above
Answer #1

. mg-T= ma -O & Txr= I x = MR xa Net quo, 5 l - from 0 » T= m (g-a) T= mcg-axr Auswer (C) =